### v1

@@ Abstract
A $(\phi,\epsilon)$-expander-decomposition of a graph $G$ (with $n$ vertices and $m$ edges) is a partition of $V$ into clusters $V_{1},\ldots,V_{k}$ with conductance $\Phi(G[V_{i}])\geq\phi$, such that there are at most $\epsilon m$ inter-cluster edges. Such a decomposition plays a crucial role in many graph algorithms. [2] gave a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n})$-expander decomposition.
In this paper we generalize this result for a broader notion of expansion. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. A standard generalization of conductance of a cut $(S,\bar{S})$ is its $\mu$-expansion $\Phi^{\mu}_{G}(S,\bar{S})=|E(S,\bar{S})|/\min\{\mu(S)),\mu(\bar{S})\}$, where $\mu(S)=\sum_{v\in S}\mu(v)$.
We present a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n}\left(\frac{\mu(V)}{m}\right))$-expander decomposition with respect to $\mu$-expansion.
@@ 1 Introduction
The conductance of a cut $(S,\bar{S})$ is $\Phi_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\operatorname{\mathbf{vol}}(S),\operatorname{\mathbf{vol}}(\bar{S}))}$, where $\operatorname{\mathbf{vol}}(S)$ is the sum of the degrees of the vertices of $S$. The conductance of a graph $G$ is the smallest conductance of a cut in $G$.
@@
A $(\phi,\epsilon)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with conductance $\Phi(G[V_{i}])\geq\phi$ such that there are at most $\epsilon m$ inter-cluster edges, where $\phi,\epsilon\geq 0$. Consider the problem of computing a $(\phi,\epsilon)$-expander decomposition for a given graph $G$ and $\phi>0$, while minimizing $\epsilon$ as a function of $\phi$. It is known that a $(\phi,\epsilon)$-expander decomposition, with $\epsilon=O(\phi\log n)$, always exists and that $\epsilon=\Theta(\phi\log n)$ is best possible [12, 1]. In a recent result, [2] achieved a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(m)$ time.
@@
The following is an overview on the result of [2] which builds on [12]. The naive way to compute an expander decomposition is to use an approximation algorithm for the sparsest cut problem. Given an $f(n)$-approximation algorithm for the problem of finding a minimum conductance cut, one can get a $(\phi,O(f(n)\cdot\phi\log n))$-expander decomposition algorithm by recursively computing approximate cuts until all components are certified as expanders. The drawback of this approach is a possible linear recursion depth (because of unbalanced cuts - one side has very small volume) which affects the running time. To solve this problem, Saranurak and Wang [12] developed a sparse cut algorithm that either returns a balanced sparse cut or an unbalanced sparse cut in which the larger side is an expander (and therefore they only need to recur on the smaller side of the cut). Their algorithm is based on the Cut-Matching Game [5].
@@
Cut-matching: The cut-matching game, played by a cut player and a matching player to construct an expander, is a well known technique to reduce the approximation task for sparsest cut to a polylogarithmic number of maximum flow problems [5]. Khandekar et al. [5] devised a strategy for the cut player that translates to an $O(\log^{2}{n})$ approximation algorithm for the sparsest cut problem. This cut player of [5] is robust in the sense that it is simple and flexible to modifications for other applications [10, 12]. A more sophisticated spectral cut player was devised by [9] which yields $O(\log{n})$ approximation. This cut player is less flexible and harder to modify.
@@
By modifying the cut-matching game and the cut player of [5], Saranrurak and Wang [12] achieved a $(\phi,\phi\log^{3}{n})$-expander decomposition in $\tilde{O}(m)$ time. In a recent result [2] were able by a careful spectral analysis, to combine the techniques of [12] and the spectral cut player of [9] to get a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(m)$ time.
@@
A useful generalization of conductance is $\mu$-expansion. Let $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. The $\mu$-expansion of a cut $(S,\bar{S})$ is $\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\mu(S),\mu(\bar{S}))}$, where $\mu(S)=\sum_{v\in S}\mu(v)$. If $\min(\mu(S),\mu(\bar{S}))=0$ then $\Phi^{\mu}_{G}(S,\bar{S})=\infty$. Note that by setting $\mu(v)=deg_{G}(v)$, we get conductance. The $\mu$-expansion of $G$ is defined as $\Phi^{\mu}(G)=\min_{S\subseteq V}\Phi^{\mu}_{G}(S,\bar{S})$. We say that $G$ is a $(\phi,\mu)$-expander if $\Phi^{\mu}(G)\geq\phi$.
@@
A $(\phi,\epsilon,\mu)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with $\mu$-expansion $\Phi^{\mu}(G[V_{i}])\geq\phi$ such that there are at most $\epsilon m$ inter-cluster edges, where $\phi,\epsilon\geq 0$ and $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$.
@@
Our Contribution: In this paper we generalize the result of [2] to the notion of $\mu$-expansion and present an $\tilde{O}(m)$ algorithm for computing a $(\phi,\phi\log^{2}{n}\frac{\mu(V)}{m})$-expander decomposition with respect to $\mu$-expansion. The generalization is quite straightforward; we derive it by modifying a simplified version of  [2] (This simplification is possible due to a recent result on fair-cut [7]). As an immediate consequence of our balanced sparse cut procedure, used in the expander decomposition algorithm (see Lemma 3.1), we get an $O(\log{n})$ approximation algorithm to the sparsest cut problem with respect to a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$.
@@
Related work: In a recent paper, Ameranis et al. [3] use a generalized notion of $\mu$-expansion, also mentioned in [8], where they allowed to consider overlapping partitions of $V$. They define a corresponding generalized version of the cut-matching game, and show how to use a cut strategy for this game to get an approximation algorithm for two generalized cut problems. They claim that one can construct a cut strategy for this measure using ideas from [8]. [^1] The details of such a cut player do not appear in [3] or [8].
@@
The structure of this paper is as follows. Section 2 contains basic definitions. In section 3 we present a version of the cut-matching game, together with a cut player and a matching player which are suitable for our expander decomposition algorithm with respect to a vertex measure $\mu$. Finally, Section 4.3 contains the balanced cut algorithm, from which we obtain an expander decomposition in a standard way which is described in Appendix 5. Appendix A contains missing proofs. Appendices B and C contain algebraic tools and probabilistic lemmas, respectively.
@@ 2 Preliminaries
Throughout the paper we work with the undirected graph $G=(V,E)$. We denote the transpose of a vector or a matrix $x$ by $x^{\prime}$. That is if $v$ is a column vector then $v^{\prime}$ is the corresponding row vector. For a vector $v\in\mathop{\mathbb{R}}^{n}$, define $\sqrt{v}$ to be vector whose coordinates are the square roots of those of $v$. Given $A\in\mathop{\mathbb{R}}^{n\times n}$, we denote by $A(i,j)$ the element at the $i$’th row and $j$’th column of $A$. We denote by $A(i,),A(,i)$ the $i$’th row and column of $A$, respectively. We think of both $A(i,)$ and $A(,i)$ as column vectors. We use the abbreviation $A(i)\mathrel{\mathop{:}}=A(i,)$ only with respect to the rows of $A$. Given a vector $v\in\mathop{\mathbb{R}}^{n}$, we denote its $i$’th element by $v(i)$. For disjoint $A,B\subseteq V$, we denote by $E_{G}(A,B)$ the set of edges connecting $A$ and $B$. We denote by $|E_{G}(A,B)|$ the number of edges in $E_{G}(A,B)$, or the sum of their weights if the graph is weighted. We sometimes omit the subscript when the graph is clear from the context. For a weighted graph $H=(V,E,w)$, we define $deg_{H}(v)=w\left(E_{H}(\{v\},V\setminus\{v\})\right)$. If $A=V\setminus B$, then we call $(A,B)$ a cut. Throughout the paper we work with the vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. For $S\subseteq V$ we let $\mu(S)=\sum_{v\in S}\mu(v)$. We define the set of terminals to be $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$. We assume the following on $\mu$.
@@
For every $v\in\mathcal{T}$ we have $1/poly(n)\leq\mu(v)\leq poly(n)$.
@@
Given a vector $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ and a collection of weighted pairs $M=\left\{(u_{i},v_{i},w_{i})\right\}_{i=1}^{m}$. We say that $M$ is a $\mu$-matching if the graph defined by $M$ satisfies $deg_{M}(v)=\mu(v)$, for every $v$.
@@
A matrix $F\in\mathop{\mathbb{R}}^{n\times n}$ is $\mu$-stochastic with respect to $\mu\in\mathop{\mathbb{R}}^{n}$ if the following two conditions hold: (1) $F\cdot\mathbbm{1}_{n}=\mu$ and (2) $\mathbbm{1}_{n}^{\prime}\cdot F=\mu^{\prime}$.
@@
Let $G=(V,E)$ and $\subset V, SS\neq\emptyset$. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. The $\mu$-expansion of the cut $(S,\bar{S})$, denoted by $\Phi^{\mu}_{G}(S,\bar{S})$, is
$\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\mu(S),\mu(\bar{S}))}.$
The $\mu$-expansion of $G$ is defined to be $\Phi^{\mu}(G)=\min_{S\subseteq V}\Phi_{G}(S,\bar{S})$.
Remark 2.6.
In the above definition we allow the denominator to be 0, in which case the $\mu$-expansion of the cut is $\infty$. Therefore, the minimization in the definition of $\Phi^{\mu}(G)$ can be taken over sets $S\subseteq V$ satisfying $\mu(S),\mu(\bar{S})>0$.
@@
$\min_{S\subseteq A}\frac{|E(S,V\setminus S)|}{\min(\mu(S),\mu(A\setminus S))}\geq\phi.$
That is, a near expander is allowed to use cut edges that go outside of $A$. Note that a subset of a near expander is also a near expander. This notion is useful in several contexts. Specifically, using “trimming” of [12], under certain conditions on $A$, one can extract an expander $B\subseteq A$ by “trimming” a part of $A$ (see Theorem 5.1).
@@
Let $G=(V,E)$ be an undirected graph. Let $F\in\mathop{\mathbb{R}}^{V\times V}_{\geq 0}$ be a matrix (not necessarily symmetric). We say that $F$ is embeddable in $G$ with congestion $c$, if there exists a multi-commodity flow $f$ in $G$, with $|V|$ commodities, one for each vertex (the vertex $v$ is the source of its commodity), such that, simultaneously for each $(u,v)\in V\times V$, $f$ routes $F(u,v)$ units of $u$’s commodity from $u$ to $v$, and the total flow on each edge is at most $c$.
If $F$ is the weighted adjacency matrix of a graph $H$ on the same vertex set $V$, we say that $H$ is embeddable in $G$ with congestion $c$ if $F$ is embeddable in $G$ with congestion cc. [^2] This definition requires to route $F(u,v)=F(v,u)$ both from $u$ to $v$ and from $v$ to $u$ if $F$ is symmetric.
@@
Let $S\subseteq A,\bar{S}=V\setminus S$ be a cut, and assume $\mu(S)\leq\mu(A\setminus S)$. In the embedding of $H$ in $G$, each edge $(u,v)\in E(H)$ corresponds to $w_{H}(u,v)$ units of flow, routed in $G$ from $u$ to $v$. Since each edge of $G$ carries at most $c$ units of flow we get that $|E_{G}(S,\bar{S})|\geq\frac{2}{c}|E_{H}(S,\bar{S})|$. [^3] Recall that by definition for every $\{u,v\}\in E(H)$ we send flow in $G$ in both directions. Since $A$ is a near $(\phi,\mu)$-expander in $H$, we get that
$\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E_{G}(S,\bar{S})|}{\mu(S)}\geq\frac{2}{c}\frac{|E_{H}(S,\bar{S})|}{\mu(S)}\geq\frac{2\phi}{c}.$

### v2

@@ Abstract
A $(\phi,\epsilon)$-expander-decomposition of a graph $G$ (with $n$ vertices and $m$ edges) is a partition of $V$ into clusters $V_{1},\ldots,V_{k}$ with conductance $\Phi(G[V_{i}])\geq\phi$, such that there are at most $\epsilon m$ inter-cluster edges. Such a decomposition plays a crucial role in many graph algorithms. [agassy2022expander] gave a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n})$-expander decomposition.
In this paper we generalize this result for a broader notion of expansion. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. A standard generalization of conductance of a cut $(S,\overline{S})$ is its $\mu$-expansion $\Phi^{\mu}_{G}(S,\overline{S})=|E(S,\overline{S})|/\min\{\mu(S),\mu(\overline{S})\}$, where $\mu(S)=\sum_{v\in S}\mu(v)$.
We present a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n}\cdot\frac{\mu(V)}{m})$-expander decomposition with respect to $\mu$-expansion. A substantial portion of the exposition is adapted from [agassy2022expander], and this work serves as a convenient reference for the generalized expander decomposition.
@@ 1 Introduction
The conductance of a cut $(S,\overline{S})$ is $\Phi_{G}(S,\overline{S})=\frac{|E(S,\overline{S})|}{\min(\operatorname{\mathbf{vol}}(S),\operatorname{\mathbf{vol}}(\overline{S}))}$, where $\operatorname{\mathbf{vol}}(S)$ is the sum of the degrees of the vertices of $S$. The conductance of a graph $G$ is the smallest conductance of a cut in $G$.
@@
A $(\phi,\epsilon)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with conductance $\Phi(G[V_{i}])\geq\phi$ such that there are at most $O(\epsilon m)$ inter-cluster edges, where $\phi,\epsilon\geq 0$. Consider the problem of computing a $(\phi,\epsilon)$-expander decomposition for a given graph $G$ and $\phi>0$, while minimizing $\epsilon$ as a function of $\phi$. It is known that a $(\phi,\epsilon)$-expander decomposition, with $\epsilon=O(\phi\log n)$, always exists and that $\epsilon=\Theta(\phi\log n)$ is best possible [saranurak2019expander, alev2017graph]. In a recent result, [agassy2022expander] achieved a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(m)$ time.
@@
The following is an overview on the result of [agassy2022expander] which builds on [saranurak2019expander]. The naive way to compute an expander decomposition is to use an approximation algorithm for the sparsest cut problem. Given an $f(n)$-approximation algorithm for the problem of finding a minimum conductance cut, one can get a $(\phi,O(f(n)\cdot\phi\log n))$-expander decomposition algorithm by recursively computing approximate cuts until all components are certified as expanders. The drawback of this approach is a possible linear recursion depth (because of unbalanced cuts - one side has very small volume) which affects the running time. To solve this problem, Saranurak and Wang [saranurak2019expander] developed a sparse cut algorithm that either returns a balanced sparse cut or an unbalanced sparse cut in which the larger side is an expander (and therefore they only need to recur on the smaller side of the cut). Their algorithm is based on the Cut-Matching Game [khandekar2009graph].
@@
Cut-matching: The cut-matching game, played by a cut player and a matching player to construct an expander, is a well known technique to reduce the approximation task for sparsest cut to a polylogarithmic number of maximum flow problems [khandekar2009graph]. Khandekar et al. [khandekar2009graph] devised a strategy for the cut player that translates to an $O(\log^{2}{n})$ approximation algorithm for the sparsest cut problem. This cut player of [khandekar2009graph] is robust, in the sense that it is simple and can be flexibly adapted for other applications [racke2014computing, saranurak2019expander]. A more sophisticated spectral cut player was devised by Orecchia et al. [orecchia2008partitioning] which yields $O(\log{n})$ approximation. This cut player is less flexible and harder to modify.
@@
By modifying the cut-matching game and the cut player of [khandekar2009graph], [saranurak2019expander] achieved a $(\phi,\phi\log^{3}{n})$-expander decomposition in $\tilde{O}(\frac{m}{\phi})$ time. In a recent result [agassy2022expander] were able by a careful spectral analysis, to combine the techniques of [saranurak2019expander] and the spectral cut player of [orecchia2008partitioning] to get a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(\frac{m}{\phi})$ time. Using the recent fair-cuts result [li2022fair], both expander decomposition algorithms [saranurak2019expander, agassy2022expander] can be implemented in $\tilde{O}(m)$ time.
@@
A useful generalization of conductance is $\mu$-expansion. Let $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. The $\mu$-expansion of a cut $(S,\overline{S})$ is $\Phi^{\mu}_{G}(S,\overline{S})=\frac{|E(S,\overline{S})|}{\min(\mu(S),\mu(\overline{S}))}$, where $\mu(S)=\sum_{v\in S}\mu(v)$. If $\min(\mu(S),\mu(\overline{S}))=0$ then $\Phi^{\mu}_{G}(S,\overline{S})=\infty$. Note that by setting $\mu(v)=\deg_{G}(v)$, we get conductance. The $\mu$-expansion of $G$ is defined as $\Phi^{\mu}(G)=\min_{S\subseteq V}\Phi^{\mu}_{G}(S,\overline{S})$. We say that $G$ is a $(\phi,\mu)$-expander if $\Phi^{\mu}(G)\geq\phi$.
@@
A $(\phi,\epsilon,\mu)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with $\mu$-expansion $\Phi^{\mu}(G[V_{i}])\geq\phi$ such that there are at most $O(\epsilon m)$ inter-cluster edges, where $\phi,\epsilon\geq 0$ and $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$.
@@
Our Contribution: In this paper we generalize the result of [agassy2022expander] to the notion of $\mu$-expansion and present an $\tilde{O}(m)$ algorithm for computing a $(\phi,\phi\log^{2}{n}\frac{\mu(V)}{m})$-expander decomposition with respect to $\mu$-expansion. The generalization is quite straightforward; we derive it by modifying a simplified version of [agassy2022expander] (this simplification is possible due to a recent result on fair-cut [LNPSsoda13]). As an immediate consequence of our balanced sparse cut procedure, used in the expander decomposition algorithm (see Lemma 3.1), we get an $O(\log{n})$ approximation algorithm to the balanced sparse cut problem with respect to a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$.
The core ideas of this paper are the same as in our previous work [agassy2022expander]. However, the generalization we pursue here alters the technical development and many internal details substantially. For this reason, rather than layering modifications onto [agassy2022expander], we provide a fresh and self-contained account of the result. Several passages are therefore intentionally adapted from [agassy2022expander] to maintain continuity and clarity of exposition.
@@
Related work: In a recent paper, Ameranis et al. [pmlr-v162-orecchia22a] use a generalized notion of $\mu$-expansion, also mentioned in [orecchia2011fast], where they allowed to consider overlapping partitions of $V$. They define a corresponding generalized version of the cut-matching game, and show how to use a cut strategy for this game to get an approximation algorithm for two generalized cut problems. They claim that one can construct a cut strategy for this measure using ideas from [orecchia2011fast]. [^1] The details of such a cut player do not appear in [pmlr-v162-orecchia22a] or [orecchia2011fast].
@@
The structure of this paper is as follows. Section 2 contains basic definitions. In section 3 we present a version of the cut-matching game, together with a cut player and a matching player which are suitable for our expander decomposition algorithm with respect to a vertex measure $\mu$. Finally, Section 4 contains the balanced cut algorithm, from which we obtain an expander decomposition in a standard way which is described in Section 5. Appendix A contains proofs that were deferred from the main text. Appendices B and C contain algebraic tools and probabilistic lemmas, respectively.
@@ 2 Preliminaries
Throughout the paper we work with the undirected graph $G=(V,E)$. Let $n\mathrel{\mathop{:}}=|V|$ and $m\mathrel{\mathop{:}}=|E|$. We denote the transpose of a vector or a matrix $x$ by $x^{\prime}$. That is if $v$ is a column vector then $v^{\prime}$ is the corresponding row vector. For a vector $v\in\mathop{\mathbb{R}}_{\geq 0}^{k}$, define $\sqrt{v}$ to be vector whose coordinates are the square roots of those of $v$. Given $A\in\mathop{\mathbb{R}}^{k\times l}$, we denote by $A(i,j)$ the element at the $i$’th row and $j$’th column of $A$. We denote by $A(i,),A(,i)$ the $i$’th row and column of $A$, respectively. We think of both $A(i,)$ and $A(,i)$ as column vectors. We use the abbreviation $A(i)\mathrel{\mathop{:}}=A(i,)$ only with respect to the rows of $A$. Given a vector $v\in\mathop{\mathbb{R}}^{k}$, we denote its $i$’th element by $v(i)$. For disjoint $A,B\subseteq V$, we denote by $E_{G}(A,B)$ the set of edges connecting $A$ and $B$. We denote by $|E_{G}(A,B)|$ the number of edges in $E_{G}(A,B)$, or the sum of their weights if the graph is weighted. We sometimes omit the subscript when the graph is clear from the context. For a weighted graph $H=(V,E,w)$, we define $\deg_{H}(v)=w\left(E_{H}(\{v\},V\setminus\{v\})\right)$. If $A=V\setminus B$, then we call $(A,B)$ a cut. Throughout the paper we work with an arbitrary vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$. [^2] We identify the set $V$ with $[n]$, and therefore we sometimes write $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. For $S\subseteq V$ we let $\mu(S)=\sum_{v\in S}\mu(v)$. We define the set of terminals to be $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$. We assume the following on $\mu$. This assumption is mainly used in the proof of Lemma 4.9.
@@
For every $v\in\mathcal{T}$ we have $\frac{1}{\operatorname{\mathbf{poly}}(n)}\leq\mu(v)\leq\operatorname{\mathbf{poly}}(n)$.
@@
Given a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ and a collection of weighted pairs $M=\left\{(u_{i},v_{i},w_{i})\right\}_{i=1}^{k}$. We say that $M$ is a $\mu$-matching if the graph defined by $M$ satisfies $\deg_{M}(v)=\mu(v)$, for every $v$.
@@
Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. A matrix $F\in\mathop{\mathbb{R}}^{n\times n}$ is called $\mu$-stochastic if the following two conditions hold: (1) $F\cdot\mathbbm{1}_{n}=\mu$ and (2) $\mathbbm{1}_{n}^{\prime}\cdot F=\mu^{\prime}$.
@@
Let $G=(V,E)$ and $\subset V, SS\neq\emptyset$. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. The $\mu$-expansion of the cut $(S,\overline{S})$ satisfying $\min(\mu(S),\mu(\overline{S}))>0$, denoted by $\Phi^{\mu}_{G}(S,\overline{S})$, is
$\Phi^{\mu}_{G}(S,\overline{S})=\frac{|E(S,\overline{S})|}{\min(\mu(S),\mu(\overline{S}))}.$
The $\mu$-expansion of $G$ is defined to be
$\Phi^{\mu}(G)=\min_{\begin{subarray}{c}S\subseteq V\\ \min(\mu(S),\mu(\overline{S}))>0\end{subarray}}\Phi_{G}(S,\overline{S})$
@@
$\min_{\begin{subarray}{c}S\subseteq A\\ \min(\mu(S),\mu(A\setminus S))>0\end{subarray}}\frac{|E(S,V\setminus S)|}{\min(\mu(S),\mu(A\setminus S))}\geq\phi.$
That is, a near expander is allowed to use cut edges that go outside of $A$. Note that a subset of a near expander is also a near expander. This notion is useful in several contexts. Specifically, using “trimming” of [saranurak2019expander], under certain conditions on $A$, one can extract an expander $B\subseteq A$ by “trimming” a part of $A$ (see Theorem 5.1).
@@
Let $G=(V,E)$ be an undirected graph. Let $F\in\mathop{\mathbb{R}}^{V\times V}_{\geq 0}$ be a matrix (not necessarily symmetric). We say that $F$ is embeddable in $G$ with congestion $c$, if there exists a multi-commodity flow $f$ in $G$, with $|V|$ commodities, one for each vertex (the vertex $v$ is the source of its commodity), such that, simultaneously for each $(u,v)\in V\times V$, $f$ routes $F(u,v)$ units of $u$’s commodity from $u$ to $v$, and the total flow on each edge is at most $c$ times the capacity of the edge.
If $F$ is the weighted adjacency matrix of a graph $H$ on the same vertex set $V$, we say that $H$ is embeddable in $G$ with congestion $c$ if $F$ is embeddable in $G$ with congestion cc. [^3] This definition requires to route $F(u,v)=F(v,u)$ units both from $u$ to $v$ and from $v$ to $u$ if $F$ is symmetric.
@@
Let $S\subseteq A,\overline{S}=V\setminus S$ be a cut, and assume $0<\mu(S)\leq\mu(A\setminus S)$. In the embedding of $H$ in $G$, each edge $(u,v)\in E(H)$ corresponds to $w_{H}(u,v)$ units of flow, routed in $G$ from $u$ to $v$. Since each edge of $G$ carries at most $c$ units of flow (times its capacity, for weighted graphs) we get that $|E_{G}(S,\overline{S})|\geq\frac{2}{c}|E_{H}(S,\overline{S})|$. [^4] Recall that by definition for every $\{u,v\}\in E(H)$ we send flow in $G$ in both directions. Since $A$ is a near $(\phi,\mu)$-expander in $H$, we get that
$\Phi^{\mu}_{G}(S,\overline{S})=\frac{|E_{G}(S,\overline{S})|}{\mu(S)}\geq\frac{2}{c}\frac{|E_{H}(S,\overline{S})|}{\mu(S)}\geq\frac{2\phi}{c}.$